amya:

At their going-out-of-business sale, Bill's Bikes sold 33 bikes and had 53 left. Which equation could be used to find x, the number of bikes the store had originally?

Madiron:

ummm

Madiron:

let me think

mackeulerick2028:

If we look at the original problem, we see that Bill's Bikes had 60 bikes left over and had sold 41 bikes, our original amount of bikes can be expressed as: Number of original bikes = Leftover Bikes + Sold Bikes Number of original bikes is the variable we are being asked to find, so: x = 60 + 41 However, as this problem is phrased, Bill's Bikes sold 41 bikes to get 60 bikes left over, thus they want you to write the equation as: x = 60 + 41 -41 -41 x - 41 = 60 A simple rearrangement of terms gives you this solution. "The Bikes we had (x) minus the bikes we sold (41) gives us the bikes left over (60)"

Madiron:

yeah what that person said

amya:

thanks

mackeulerick2028:

np